Azree an accountant, working for a foreign consultant firm and earning RM 78,000 per year is contemplating giving up his job and set up his own tax consultant firm. He estimates that renting an office would cost RM 780 per month, hiring a secretary with salary RM 1,500 per month and purchasing for required supplies would cost him RM 10,000 per annum. He estimated that his total revenues for the year would be RM 120,000. Calculate  explicit cost, implicit cost, accounting profit & economic profit.  

   


1
Expert's answer
2021-04-30T10:47:39-0400

Explicit cost is the sum of all costs :10,000+(780+1,500)×12=37,360. Implicit cost includes also the lost salary: 37,360+78,000=115,360. Accounting profit is 120,000-37,360=82,640; economic profit is 120,000-115,360=4640